OpenStudy (anonymous):

Find the first four terms of the multiplication pattern given by the formula. 10 • 4n – 1

OpenStudy (anonymous):

@skay

OpenStudy (anonymous):

@ladydeadpool

OpenStudy (anonymous):

@qweqwe123123123123111

OpenStudy (anonymous):

@InsanelyChaotic

OpenStudy (anonymous):

@TSwizzle

OpenStudy (anonymous):

@Falco276

OpenStudy (anonymous):

@Reppinthe80

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@ccooll

OpenStudy (anonymous):

@julian25

OpenStudy (anonymous):

@ginlynn17

OpenStudy (anonymous):

"n" is a number. What are the first 4 numbers? Plug them into your equation and you have your answer.

OpenStudy (anonymous):

Put in 1, 2, 3, then 4 for n 10•4 - 1 = 39 10•8 - 1 = 79 10•12 - 1 = 119 10•16 - 1 = 159

OpenStudy (anonymous):

thx so much

OpenStudy (anonymous):

Of Course !

OpenStudy (falco276):

Congrats! You got the answer! :D